### v1

# VIRUS-NeRF - Vision, InfraRed and UltraSonic based Neural Radiance Fields

Abstract
Autonomous mobile robots are an increasingly integral part of modern factory and warehouse operations. Obstacle detection, avoidance and path planning are critical safety-relevant tasks, which are often solved using expensive LiDAR sensors and depth cameras. We propose to use cost-effective low-resolution ranging sensors, such as ultrasonic and infrared time-of-flight sensors by developing VIRUS-NeRF - Vision, InfraRed, and UltraSonic based Neural Radiance Fields.
Building upon Instant Neural Graphics Primitives with a Multiresolution Hash Encoding (Instant-NGP), VIRUS-NeRF incorporates depth measurements from ultrasonic and infrared sensors and utilizes them to update the occupancy grid used for ray marching. Experimental evaluation in 2D demonstrates that VIRUS-NeRF achieves comparable mapping performance to LiDAR point clouds regarding coverage. Notably, in small environments, its accuracy aligns with that of LiDAR measurements, while in larger ones, it is bounded by the utilized ultrasonic sensors. An in-depth ablation study reveals that adding ultrasonic and infrared sensors is highly effective when dealing with sparse data and low view variation. Further, the proposed occupancy grid of VIRUS-NeRF improves the mapping capabilities and increases the training speed by 46% compared to Instant-NGP. Overall, VIRUS-NeRF presents a promising approach for cost-effective local mapping in mobile robotics, with potential applications in safety and navigation tasks. The code can be found at https://github.com/ethz-asl/virus_nerf.
I INTRODUCTION
As automation advances, the demand for mobile robots is on the rise. In the realm of factories and warehouses, Atonomous Mobile Robots exhibit remarkable flexibility and perform tasks with greater intelligence compared to traditional Automated Guided Vehicles. Typically, AMRs operate in semi-dynamic environments alongside human workers. The robot must effectively perceive its surroundings to facilitate smooth navigation and obstacle avoidance. Operating within the same workspace as humans requires a safe mapping algorithm, focusing particularly on the proximity of the robot (e.g. see Fig. 1). In this work, perception is studied in the simplified case of static environments and the mapping is evaluated in 2D space.
In industry, safety-critical tasks for AMRs, such as collision avoidance, often rely on costly sensors, such as 3D Light Detection And Ranging (LiDAR) sensors. Typically, local and instantaneous measurements are utilized being more robust and allowing faster scanning cycles than when using global mapping. This study seeks to employ cost-efficient sensors while maintaining local mapping capabilities similar to more expensive setups. For example, Ultrasonic Sensors are widely adapted low-cost ranging sensors in the car industry [1] and many of the early mapping algorithms in mobile robotics utilize them [2, 3, 4]. However, their low angular resolution combined with traditional fusion methods lead to sub-optimal outcomes. Similarly, other cheap time-of-flight sensors, e.g. Infrared Sensors, result in limited mapping performances due to their sparse measurements and reduced range. Contemporary approaches frequently rely on more advanced sensors like LiDARs or depth cameras. Despite their impressive performance, these setups incur substantial costs, making them less accessible for widespread adoption. Recent advancements in stereo vision [5, 6] and Monocular Depth Estimation (MDE) [7, 8, 9] show significant progress in camera-based mapping solutions. However, the lack of robustness, high computational requirements and the intrinsic scale ambiguity of monocular cameras remain open research challenges. Sensor fusion may address some of these drawbacks, e.g. depth completion [10, 11, 12]. Nevertheless, most of these techniques return to using costly LiDAR or RGB-D sensors.
This study suggests using the framework of Neural Radiance Fields to fuse color images with range measurements and to learn an implicit scene representation. In the context of mobile robotics, NeRFs incur two major drawbacks: First, they converge slowly, which makes them problematic for real-time usage, and second, they require dense data with a high view variation of the environment [13, 14]. More recent works address the convergence speed by proposing various improvements [13, 11]. Most mobile robots, especially in warehouse and factory environments, are constrained to move along pre-defined trajectories on a 2D plane, severely limiting their viewpoint variability. Therefore, we developed VIRUS-NeRF - a Vision, InfraRed and UltraSonic based Neural Radiance Fields. VIRUS-NeRF is based on Instant-NGP. Similar to other works using LiDARs [15, 16] or depth cameras [17, 18], VIRUS-NeRF complements the image-based training by depth measurements. Notably, VIRUS-NeRF is the first NeRF algorithm utilizing low-cost depth sensors, i.e. USSs and IRSs. The contributions of this work are the following:
A novel real-time, NeRF-based sensor fusion method for integrating low-cost and low-resolution USSs and IRSs with RGB cameras. The low-cost sensors provide depth supervision, to the normally purely image-based training of NeRFs, resulting in more accurate and robust reconstructions of the environment.
Improvements to the occupancy grid of Instant-NGP using a probabilistic Bayesian formulation, which permits direct occupancy updates by depth measurements.
An in-depth ablation study comparing VIRUS-NeRF to Instant-NGP, analyzing the contribution of the depth supervision and the improved occupancy grid separately, and studying different isolated sensor modalities.
II Related Work
II-A Occupancy Grids
One of the first mapping techniques is the occupancy grid, which is also utilized in Instant-NGP, the base model of VIRUS-NeRF (see section II-C). Occupancy grids are probabilistic maps describing the occupancy state of a discretized environment [2]. Many of the early publications explicitly use low-cost sensors, e.g. USSs [2, 19, 20, 21]. The most established implementation is based on the Bayesian Updating Rule to integrate new measurements into the map [3]. VIRUS-NeRF makes use of the Bayesian occupancy grid (see section III-B) which considers consecutive measurements and neighboring cells independently to reduce the updating complexity. These strong assumptions may be dropped by addressing the correlation between successive samples [20] or between neighboring cells [21]. Multiple works extend occupancy grids to handle dynamic objects using Bayesian Occupancy Filters [22], Particle Filters [23] or Markov Chains [24, 25].
II-B Depth Completion
Monocular cameras provide extremely rich information and are available for moderate prices, therefore being extensively used in mobile robotics. VIRUS-NeRF employs monocular cameras in the framework of NeRFs. However, there exist some viable methods based on deep learning: For example, MDE predicts the pixel-wise depth by using color images [8]. RGB images have no inherent depth information and scale ambiguity limits the accuracy of MDE. Additionally, most MDE algorithms do not estimate the uncertainty of their predictions [26], which makes sensor fusion difficult. Contrary, depth completion tries to leverage the early fusion of RGB images with depth measurements. Most approaches are based on deep learning and complement the camera with a LiDAR sensor [11]. As an alternative to LiDAR point clouds, radar scans can be used for depth completion [27, 28, 29]. These models are relatively new because of the availability of large-scale open-source datasets containing radar data [30] and the improved resolution of radar sensors in the past few years [28]. Existing methods developed for LiDAR do not transfer well to radars, due to the sparsity and high noise levels of the data, as well as the smaller vertical Field of View (FoV) [27, 29]. To the best of our knowledge, there is no published work on utilizing USSs or IRSs for depth completion and there does not exist any large-scale open-source dataset containing such low-cost sensors.
II-C Neural Radiance Fields
In 2020, Mildenhall et al. introduce NeRFs [31]. NeRFs use Multi-Layer Perceptrons (MLP) to learn the geometry and the lighting of one particular three-dimensional static scene. During inference, the model can render a new view from any position and viewing direction. The required data is composed of images and their respective camera poses.
The original NeRF implementation [31] has some important limitations: The training of one scene takes a long time (up to 12h on a GPU [13]). In addition, NeRFs are designed for small scenes, i.e. single objects or small rooms, and struggle with unbounded environments [16]. Dense data having a high view variation is required for training [13, 14] and surfaces can be rugged due to a lack of geometrical constraints [32]. Many of these drawbacks are addressed in subsequent publications, as presented below.
Instant-NGP reduces the training time from several hours to a few minutes while achieving a similar accuracy [33]. Instead of a sinusoidal encoding like in the original implementation, Instant-NGP uses a multi-resolution hash encoding. In addition, Instant-NGP proposes to use a occupancy grid, making the ray marching more efficient. The grid is updated by a heuristic rule using density predictions and thresholded by a fixed value to distinguish occupied space, where points are sampled, from unoccupied areas, which are skipped. The current state-of-the-art in terms of surface reconstruction (including large outdoor scenes) is Neuralangelo [34]. Neuralangelo is based on Instant-NGP optimizing the hash grid with a coarse-to-fine approach. Similar to NeuS [32], Neuralangelo uses signed distance functions and is improved by adding smoothing constraints.
Urban Radiance Fields [15], CLONeR [16], iMAP [17] and NICE-SLAM [18] reduce the demand for dense data by adding depth supervision. Urban Radiance Fields and CLONeR use LiDAR point clouds in addition to color images. CLONeR separates the occupancy and the color into two MLPs where the occupancy MLP is trained with LiDAR point clouds and the color MLP with images. iMAP [17] and NICE-SLAM [18] are Simultaneous Localization and Mapping (SLAM) algorithms and employ depth supervision from an RGB-D camera. More recent implementations, e.g. NeRF-SLAM [35] and Orbeez-SLAM [36], perform SLAM utilizing only monocular cameras by separating pose estimation from neural scene representation and leveraging visual odometry. For example, NeRF-SLAM uses Droid-SLAM [37] as a tracking module and Instant-NGP for scene representation.
NeRF is the preferred approach in this research because of its implicit sensor fusion of color images and range measurements (see section III-A). Meanwhile, it is a mapping framework having the following advantages: NeRFs are continuous and not discrete which allows in general a higher resolution. Implicit scene representations are more memory-efficient than explicit ones. For example, the smallest room with a volume of about is represented by less than in our experiments. Comparably, a 3D occupancy grid with resolution is more than 16 times larger. Besides occupancy, NeRFs learn color and lighting properties which could be used for further tasks, e.g. object classification or scene segmentation.
III VIRUS-NeRF
VIRUS-NeRF is based on Instant-NGP considering its fast convergence speed and its wide adaption [34, 35]. We propose two improvements on top of the base model: First, similarly to other works [15, 16, 17, 18], depth supervision is added to the color-based training, reducing the demand of dense data with a high view variation. However, instead of using expensive LiDARs or depth cameras, VIRUS-NeRF is based on low-cost USSs and IRSs (see chapter III-A). Second, the occupancy grid of Instant-NGP is updated by the depth measurements (see chapter III-B), making ray marching more efficient and improving the results.
III-A Depth Supervision
III-A1 Color Rendering
In general, NeRFs are trained as follows: During ray marching, a ray is traced in the viewing direction of every pixel. pairs of positions and directions are sampled along this ray. These samples are the input to a MLP that is only a few layers deep. The network predicts the color () and density () of each sample. Then, through volume rendering, the actual color of the ray () is estimated:
where is the distance between adjacent samples and is the light transmittance:
Finally, the squared error between all estimated ray colors () and the corresponding pixels () is calculated for one batch of pixels. This loss () is used for back-propagation:
III-A2 Depth Rendering
As shown in iMAP [17], the depth of a pixel can be estimated during volume rendering:
where is the depth of sample , is the distance between adjacent samples and is the light transmittance (see equation 2). The depth rendering described in equation 4 is equivalent to the color rendering of equation 1, except that the predicted depths are summed up instead of the colors .
IRS measurements are considered to be point-like, i.e. one measurement corresponds to one or few camera pixels in a close neighborhood. Analogue to the colors, the depth loss is the squared error between all estimated depths and the depth measurements :
USSs have a wide opening angle and the exact location of the object reflecting the sound wave is unknown. This prohibits applying the same depth loss as for the IRS. However, neglecting complete absorption and specular reflections of sound waves, an error can be calculated for all predictions that are closer than the measurement:
where corresponds to the accuracy of the USS. The total loss is given by the following equation:
III-A3 Rendering Bias
Volume rendering is a weighted sum of distances with weights (see equation 4). Let’s assume that the densities are described by a positive symmetric function around the surface of an object (e.g. normal distribution: center = predicted surface location, std = uncertainty). Then, the second part of the weights adopts the same symmetry as the densities. The light transmittance is a monotonically decreasing function (see equation 2). Hence, the weighting is on average larger for samples before the surface of the object than afterwards and therefore, the depth is underestimated systematically. However, the NeRF is not tied to model symmetric density functions and the bias can be absorbed into the neural network. Moreover, a few samples of high density may determine the depth estimation completely because the transmittance decreases exponentially and converges fast to zero.
III-B Occupancy Grid
III-B1 Bayesian Updating Rule
The occupancy grid of VIRUS-NeRF is based on a dual updating mechanism using the NeRF predictions similar to Instant-NGP and additionally the depth measurements. The key difference to Instant-NGP is that the occupancy grid of VIRUS-NeRF contains values in instead of . This allows for a probabilistic formulation and the use of the Bayesian Updating Rule [3], which updates the probability of a cell being occupied or empty based on the probability of making a measurement :
where is the posterior probability given measurements and is the occupancy grid before integrating . can be assumed to be equal to 0.5, or it can be derived from initial information about the map.
III-B2 Depth-Update
The probability of a depth sensor making a measurement is given by the Multiple Target Model from the MURIEL method [20] proving to be the best real-time occupancy grid in a benchmark comparing different algorithms [38]. Early-stage testing shows that the low angular resolution of USSs prevent the occupancy grid from refining. Therefore, the Depth-Update is done uniquely based on IRS measurements.
III-B3 NeRF-Update
The NeRF-Update uses no conventional measurement. However, the term can be thought of as the probability of the NeRF predicting the density given that a cell is occupied. The MLP outputs density predictions between zero and infinity. To be able to use the Bayesian Updating Rule, we introduce the projection of the density from to :
where is the predicted density by the NeRF, is the slope of the mapping function and is the density threshold. If , then the projection becomes a step function. If , then the occupancy probability is larger than and vice versa. The weights of the density MLP are randomly initialized in leading to small values at the beginning of the training. If is fixed, then would vanish in the first few cycles. Therefore, is defined as a function of as follows: where is the maximum density threshold and is the batch size. For all tested models, becomes constant after a few training steps, s.t. .
IV Experiments
IV-A Implementation
In this work, the Taichi implementation111https://github.com/Linyou/taichi-ngp-renderer [39] is employed because the original implementation is written in Cuda and therefore does not run on a CPU which is required for development.
IV-B Dataset
IV-B1 Environment
The dataset is collected by a mobile robot in two environments: Office (room with tables, chairs and cupboards, ) and Common Area (room with tables, couches and kitchen corner, ). The scenes are captured quasi-statically containing only minimal movements, e.g. a person writing on a keyboard. The robot explores the environments on 2D trajectories and makes the measurements with sensors having an overlapping FoV. This leads to a relatively low view variation of the scene.
IV-B2 Hardware
The Super Mega Bot (SMB) is used - a differential drive robot designed by Inspector Bots222https://www.inspectorbots.com/ (see Fig. 1(a)). On top of the SMB, a RoboSense RS-LiDAR-16 and two sensor stacks are fixed. The sensor stack mounts the USS DFRobot URM37333https://wiki.dfrobot.com/URM37_V5.0_Ultrasonic_Sensor_SKU_SEN0001_, the IRS STMicroelectronics VL53L5CX444https://www.st.com/en/imaging-and-photonics-solutions/vl53l5cx.html and the camera Intel RealSense D455555https://www.intelrealsense.com/depth-camera-d455/ (see Fig. 1(b)). They are designed such that all sensors of one stack are oriented in the same direction and are as close as possible. The IRS is a time-of-flight sensor measuring an array of 8x8 pixels. The stacks are approximately apart and point in to the driving direction. The configuration of the cameras having an overlapping FoV is chosen to use the calibration tool Kalibr [40] which estimates the intrinsic and extrinsic camera parameters simultaneously. The calibration concerning the LiDAR is done using the Camera-LiDAR Calibration - V 2.0 [41]. The influence of an orientation error between the IRS and the camera is tested by simulating IRS measurements and adding an artificial angular error. The Nearest Neighbour Distance (NND) is not affected by an error up to and therefore, the IRS and USS calibration obtained from Computer Aided Design (CAD) is sufficient. The sensors are not hardware-synchronized but recorded on the same clock and the temporally closest measurements are assigned during post-processing for each sensor stack.
IV-B3 Localisation
IV-C Metrics
Traditionally, safety-relevant navigational tasks are done using instantaneous depth measurements. Therefore, VIRUS-NeRF is compared to momentary scans of USSs, IRSs and LiDARs. For every scene, a global map is created by projecting the LiDAR point clouds to a grid in world coordinates using the optimized poses. This global map has a cube size and is thresholded at a minimum of two points per voxel to reduce noise. For every test point, the Ground Truth (GT) consists of a 2D depth scan at the height of the cameras within the global map. The depth predictions are not compared directly to the global map to mitigate the erroneous association of objects. For example, a depth prediction can be too far away, but another object is present at this location in the global map. This would lead to a small error even though the prediction is false. For VIRUS-NeRF, the predictions consist of a 2D depth scan at the height of the cameras, which is created by volume rendering (see equation 4). For the depth sensors, the predictions are obtained by collapsing the measurements to a 2D representation in a vertical range of above and below the camera height.
To compare the scans, the NND is calculated which is less sensitive to orientation errors than the Root Mean Square Error (RMSE). The NND can be calculated in two directions: The distance from every prediction point to the closest GT point is a measure of accuracy. The inverse direction describes the coverage of the GT by the prediction. The NND is given by the mean of all points and an inlier-outlier metric. Inliers are defined as points where the NND is less than . The assessment of accuracy, coverage, and inlier ratio aligns with the evaluation criteria employed in iMAP [17] and NICE-SLAM [18], wherein the mapping coverage is denoted as completion and the proportion of inliers as completion ratio, using a threshold of instead of . Finally, all metrics are determined for three zones defined by the GT depth. The zones roughly represent different applications: The first zone () concerns safety applications, the second one () tasks like obstacle detection and the third one () path planning.
IV-D Mapping
IV-D1 Results
The average statistics for all test points and over 10 runs are summarized in Fig. 3 for the office and the common area environment. While the amplitude of the metric depends on the particular environment, the tendency is everywhere likewise: The USS has the worst accuracy of all sensors. Up to zone 2 (), its coverage is close to the one of the LiDAR due to its large opening angle. However, in the third zone (), the coverage of the USS worsens significantly. The IRS achieves the best accuracy while having the worst coverage due to sparse measurements. The inferior accuracy of the LiDAR sensor compared to the IRS cannot be explained by the higher range of the LiDAR sensor because it is also present in zones 1 and 2. LiDARs retain the best coverage in the third zone ().
VIRUS-NeRF scores a comparable coverage than the LiDAR. The accuracy of the NeRF depends on the scene: For smaller scenes, e.g. the office, it is slightly better than LiDARs but for larger ones, e.g. the common area, it exhibits performance akin to USS. The outliers () can be separated into predictions that are too close to the robot or too far away relative to the GT. Analyzing this distinction for the coverage in zone three () shows that the largest part of all predictions is too close with approximately three-quarters of all outliers in the common area and 90% in the office. This trend is also visible in Fig. 4: VIRUS-NeRF tends to underestimate the distance to objects, leading to false-positive predictions.
IV-D2 Discussion
The results reflect the theoretical advantages and weaknesses of the sensors: The accuracy of the USS is limited by the poor angular resolution. The acceptable coverage at short range is in line with common USS applications where coverage is more important than accuracy, e.g. for car parking assistants [1]. The IRSs have an impressive accuracy, being at least two orders of magnitude cheaper than the LiDAR sensor. However, the LiDAR sensor seems to be the best trade-off between accuracy and coverage.
The reduced accuracy of VIRUS-NeRF in the common area compared to the office may be explained by the IRS having a short range of . In the office, 35.7% of all IRS measurements are valid compared to only 10.3% in the common area where the objects are more spread out. However, for safety-relevant tasks, e.g. obstacle detection, the coverage is more important than accuracy and this one remains comparable to LiDAR point clouds in all environments. VIRUS-NeRF predicts more outliers too close leading to false-positive predictions. When visualizing the results as in Fig. 4, the estimation in front of the robot is usually accurate and the hallucinations are either sideways to the trajectory (e.g. in Fig. 3(b)) or further away from the current position of the robot (e.g. in Fig. 3(b)). Lateral to the path only a few measurements are taken causing erroneous predictions. This could be addressed by adding sensors pointing in these directions. Distant hallucinations are most likely caused by the volume rendering, which is biased towards underestimating the depth as explained in chapter III-A3. If the robot moves closer to a particular region, then fewer dispensable samples influence the volume rendering and the false-positive predictions disappear, e.g. compare in Fig. 3(a) and 3(b). In contrast to path planning, hallucinations are not as critical for safety-relevant tasks, e.g. collision avoidance, especially if they vanish when moving closer.
IV-E Ablation Study
IV-E1 Results
In the ablation study, VIRUS-NeRF is compared to Instant-NGP. Additionally, the contribution of the depth supervision and the improved occupancy grid are studied, and different sensor modalities are analyzed. VIRUS-NeRF is developed and the hyper-parameters are fine-tuned in the office environment and the results of the common area show if the model generalizes well. The mean and inlier percentage are shown in table I. For the common area, the RGB-D camera outperforms all other sensor constellations because its depth images are denser and more accurate than low-cost alternatives (compare Fig. 4(a) and 4(b)). The second-best results are achieved by using VIRUS-NeRF. When removing either the USS or the IRS, the performance drops. The original Instant-NGP implementation is significantly worse in all metrics. When adding the depth losses and still using the occupancy grid of Instant-NGP, the results improve but do not reach the same level as VIRUS-NeRF. The same ablation study is repeated for the office environment: The results are similar, with the exception that omitting pose optimization is less severe or even better in terms of inlier metrics than in the common area, as smaller scenes are less affected by odometry drift. The occupancy grid of Instant-NGP has slightly better mean coverage than VIRUS-NeRF in the office.
Occ. Grid | Mean [m] | Inliers [%] | |||
/ Sensors | Scene |
Acc. |
Cov. |
Acc. |
Cov. |
Instant-NGP [33] |
C |
0.712 |
1.287 |
0.056 |
0.059 |
CAM |
O | 0.281 | 0.497 | 0.221 | 0.201 |
Instant-NGP |
C |
0.509 |
0.501 |
0.131 |
0.22 |
CAM, USS, IRS |
O | 0.164 |
[…]
0.497 | 0.506 |
VIRUS-NeRF |
C |
0.704 |
1.412 |
0.052 |
0.056 |
CAM |
O | 0.277 | 0.476 | 0.231 | 0.215 |
VIRUS-NeRF |
C |
0.49 |
0.568 |
0.146 |
0.146 |
CAM, USS |
O | 0.262 | 0.349 | 0.212 | 0.176 |
VIRUS-NeRF |
C |
0.625 |
0.643 |
0.097 |
0.169 |
CAM, IRS |
O | 0.23 | 0.374 | 0.415 | 0.46 |
VIRUS-NeRF |
C |
[…]
RGB-D |
O |
[…]
VIRUS-NeRF |
C |
0.728 |
0.922 |
0.113 |
0.186 |
Poses not opt. |
O | 0.168 | 0.238 |
[…]
VIRUS-NeRF |
C |
[…]
CAM, USS, IRS | O |
[…]
0.237 | 0.528 | 0.531 |
IV-E2 Discussion
It may be surprising that mapping without depth supervision produces poor results while standard NeRF [31] as well as Instant-NGP [33] are solely based on images showing good results in their respective studies, and other variants can represent large environments [34]. However, the dataset of this project reflects a real robot trajectory and therefore sparse measurements and a low view diversity compared to most other datasets. For example, iMAP [17], NICE-SLAM [18] and NeRF-SLAM [35] are assessed on the indoor scenes of the Replica dataset [44] sampling twice as many images along a random trajectory. Therefore, the images of Replica are characterized by significantly greater variations in viewing perspectives.
IV-F Training Speed
All training is done on a Nvidia Titan Xp GPU. The absolute speed of the presented results is expected to be approximately 20% faster if using the Cuda implementation of Instant-NGP instead of the Taichi implementation utilized in this study [39]. Fig. 5(a) shows the average NND over 10 runs for the office environment during training. Using the entire dataset, VIRUS-NeRF converges after around while being significantly faster than Instant-NGP as explained in the following.
IV-F1 Occupancy Grid
On average, VIRUS-NeRF makes 11.64 training steps per second, in contrast to 7.96 when using the grid of Instant-NGP. This is a speed-up of 46% and can be explained by two factors: Instant-NGP samples during the first 256 training steps all grid cells and afterwards a quarter, which is significantly more than 1024 samples used in VIRUS-NeRF. Additionally, the occupancy grid of VIRUS-NeRF relies partially on probabilistic sensor models (i.e. Depth-Update, see section III-B2) instead of inferring with the NeRF network which is computationally less expensive. Accelerating the training process with the occupancy grid of VIRUS-NeRF could be very interesting for any real-time application based on depth supervised Instant-NGP, e.g. NeRF-SLAM [35].
IV-F2 Offline vs. Online
Up to here, all results are generated in offline mode where all the data is available from the beginning of the training. However, in most real-world applications the mapping algorithm should already be functional before all data is collected. Online operation is using uniquely the measurements that would be available up to this time point and data playback is performed in real-time. In this case, the training lasts for the duration of the experiment and is evaluated exclusively on already visited poses. For online operation, the NND converges after which is significantly slower than after during offline training (see Fig. 6). Similarly, the Peak Signal to Noise Ratio (PSNR) takes online much longer to converge. The final NND for online learning is compared to in the offline case. The common area has similar results. However, the metrics worsen again towards the end of the experiment.
The computational speed and the training data are possible limitations of the convergence speed. It seems that the available data is more important compared to the computational power because, despite making more training steps, the online algorithm converges much slower than the offline one. In the online training of the office environment, the NND starts to converge after half of the training process being approximately the moment when the robot turns around and drives back towards its starting position. Similarly, the metrics degrade in the common area when the robot makes a sharp turn and starts moving to an unseen part of the scene. These observations and the general performance difference between offline and online training suggest that a higher variety of viewpoints would improve the convergence significantly. This is in line with most NeRF algorithms that rely on a plurality of viewing angles [13, 14] and could be addressed by adding more sensors to the robot pointing sideways and backwards.
V Conclusion
This study presents VIRUS-NeRF - Vision, InfraRed and UltraSonic based Neural Radiance Fields for local mapping. VIRUS-NeRF utilizes low-cost USSs and IRSs by adapting the depth supervision to sensors having a poor angular resolution. Additionally, the algorithm uses the sensors to update the occupancy grid introduced in Instant-NGP [33] which is used for ray marching. Two datasets are collected to evaluate the algorithm and to compare it to instantaneous scans of USSs, IRSs and LiDARs in 2D. The results show that VIRUS-NeRF has comparable coverage to LiDARs and is much better than USSs and IRSs. The accuracy of VIRUS-NeRF depends on the environment: For smaller scenes (office), it is slightly better than LiDARs but for larger ones, (common area), it exhibits accuracy akin to USS. Larger environments could be addressed by taking a IRS having a wider range.
The ablation study shows that the base model Instant-NGP has substantially worse results compared to VIRUS-NeRF. Adding the USS and the IRS to the RGB image-based training is very effective, even though the low-cost sensors have a poor angular resolution and make sparse measurements respectively. Only the more expensive RGB-D camera outperforms this sensor configuration. The occupancy grid of VIRUS-NeRF improves the metrics compared to the one of Instant-NGP and it makes the algorithm 46% faster. Generally, the convergence speed and accuracy could be improved by adding more sensors and taking measurements with a higher view variation. This research shows that VIRUS-NeRF is an effective method for local mapping based on a low-cost sensor setup.
- AGV
- Automated Guided Vehicle
- AMR
- Atonomous Mobile Robot
- ASL
- Autonomous Systems Lab
- FoV
- Field of View
- GT
- Ground Truth
- IMU
- Inertial Measurement Unit
- Instant-NGP
- Instant Neural Graphics Primitives with a Multiresolution Hash Encoding
- ICP
- Iterative Closest Point
- IRS
- Infrared Sensor
- LiDAR
- Light Detection And Ranging
- MDE
- Monocular Depth Estimation
- MLP
- Multi-Layer Perceptrons
- NeRF
- Neural Radiance Field
- NND
- Nearest Neighbour Distance
- PSNR
- Peak Signal to Noise Ratio
- RMSE
- Root Mean Square Error
- RGB-D
- Red Green Blue Depth
- ROS
- Robot Operating System
- SLAM
- Simultaneous Localization and Mapping
- SMB
- Super Mega Bot
- USS
- Ultrasonic Sensor
- CAD
- Computer Aided Design

### v2

@@
 The Super Mega Bot (SMB) is used - a differential drive robot designed by Inspector Bots222https://www.inspectorbots.com/ (see Fig.
-1(a)).
+2(a)).
 On top of the SMB, a RoboSense RS-LiDAR-16 and two sensor stacks are fixed.
 The sensor stack mounts the USS DFRobot URM37333https://wiki.dfrobot.com/URM37_V5.0_Ultrasonic_Sensor_SKU_SEN0001_, the IRS STMicroelectronics VL53L5CX444https://www.st.com/en/imaging-and-photonics-solutions/vl53l5cx.html and the camera Intel RealSense D455555https://www.intelrealsense.com/depth-camera-d455/ (see Fig.
-1(b)).
+2(b)).
 They are designed such that all sensors of one stack are oriented in the same direction and are as close as possible.
@@
 in Fig.
-3(b)) or further away from the current position of the robot (e.g.
+4(b)) or further away from the current position of the robot (e.g.
 in Fig.
-3(b)).
+4(b)).
 Lateral to the path only a few measurements are taken causing erroneous predictions.
@@
 compare in Fig.
-3(a) and 3(b).
+4(a) and 4(b).
 In contrast to path planning, hallucinations are not as critical for safety-relevant tasks, e.g.
@@
 For the common area, the RGB-D camera outperforms all other sensor constellations because its depth images are denser and more accurate than low-cost alternatives (compare Fig.
-4(a) and 4(b)).
+5(a) and 5(b)).
 The second-best results are achieved by using VIRUS-NeRF.
@@
 Occ. Grid | Mean [m] | Inliers [%] | |||
-/ Sensors | Scene |
-Acc. |
-Cov. |
-Acc. |
-Cov. |
-Instant-NGP [33] |
-C |
-0.712 |
-1.287 |
-0.056 |
-0.059 |
-CAM |
-O | 0.281 | 0.497 | 0.221 | 0.201 |
-Instant-NGP |
-C |
-0.509 |
-0.501 |
-0.131 |
-0.22 |
-CAM, USS, IRS |
-O | 0.164 |
+/ Sensors | Scene | Acc. | Cov. | Acc. | Cov. |
+Instant-NGP [33] | C | 0.712 | 1.287 | 0.056 | 0.059 |
+CAM | O | 0.281 | 0.497 | 0.221 | 0.201 |
+Instant-NGP | C | 0.509 | 0.501 | 0.131 | 0.22 |
 […]
-0.497 | 0.506 |
-VIRUS-NeRF |
-C |
-0.704 |
-1.412 |
-0.052 |
-0.056 |
-CAM |
-O | 0.277 | 0.476 | 0.231 | 0.215 |
-VIRUS-NeRF |
-C |
-0.49 |
-0.568 |
-0.146 |
-0.146 |
-CAM, USS |
-O | 0.262 | 0.349 | 0.212 | 0.176 |
-VIRUS-NeRF |
-C |
-0.625 |
-0.643 |
-0.097 |
-0.169 |
-CAM, IRS |
-O | 0.23 | 0.374 | 0.415 | 0.46 |
-VIRUS-NeRF |
-C |
+VIRUS-NeRF | C | 0.704 | 1.412 | 0.052 | 0.056 |
+CAM | O | 0.277 | 0.476 | 0.231 | 0.215 |
+VIRUS-NeRF | C | 0.49 | 0.568 | 0.146 | 0.146 |
+CAM, USS | O | 0.262 | 0.349 | 0.212 | 0.176 |
+VIRUS-NeRF | C | 0.625 | 0.643 | 0.097 | 0.169 |
+CAM, IRS | O | 0.23 | 0.374 | 0.415 | 0.46 |
+[…]
+[…]
+VIRUS-NeRF | C | 0.728 | 0.922 | 0.113 | 0.186 |
 […]
@@
 […]
-[…]
-RGB-D |
-O |
-[…]
-[…]
-[…]
-[…]
-VIRUS-NeRF |
-C |
-0.728 |
-0.922 |
-0.113 |
-0.186 |
-Poses not opt. |
-O | 0.168 | 0.238 |
-[…]
-[…]
-VIRUS-NeRF |
-C |
-[…]
-[…]
-[…]
-[…]
-CAM, USS, IRS | O |
-[…]
-0.237 | 0.528 | 0.531 |
 IV-E2 Discussion
@@
 Therefore, the images of Replica are characterized by significantly greater variations in viewing perspectives.
-IV-F Training Speed
+IV-F Training
 All training is done on a Nvidia Titan Xp GPU.
-The absolute speed of the presented results is expected to be approximately 20% faster if using the Cuda implementation of Instant-NGP instead of the Taichi implementation utilized in this study [39].
+The algorithm is expected to be approximately 20% faster if using the Cuda implementation of Instant-NGP instead of the Taichi implementation utilized in this study [39].
+IV-F1 Speed
 Fig.
-5(a) shows the average NND over 10 runs for the office environment during training.
-Using the entire dataset, VIRUS-NeRF converges after around while being significantly faster than Instant-NGP as explained in the following.
-IV-F1 Occupancy Grid
+6(a) shows the average NND over 10 runs for the office environment during training.
+Using the entire dataset, VIRUS-NeRF converges after around .
 On average, VIRUS-NeRF makes 11.64 training steps per second, in contrast to 7.96 when using the grid of Instant-NGP.
@@
 NeRF-SLAM [35].
+To further increase the convergence speed, the same forward pass could be used for the NeRF-Update of the occupancy grid as for the training of the MLP instead of treating these tasks separately.
+Furthermore, the density MLP could be partially trained without the color MLP leveraging the existence of depth measurements.
 IV-F2 Offline vs. Online